### v1

@@ 1 Introduction
This exploration has prompted a controversial discussion, the essence of which we highlight with a few selected examples. In 2011, Qian and Eberly suggested a new interpretation of light polarization based on “non-quantum entanglement," stating that “polarization is a characterization of the correlation between the vector nature and the statistical nature of the light field." Subsequent elaborations on the emerging links between quantum and classical optics [7, 8] led Eberly et al. to a strong statement: “… entanglement is a vector space property, present in any theory with a vector space framework. There is no distinction between quantum and classical entanglements, as such" [8]. This assertion has elicited equally strong opposing statements. Karimi and Boyd, for instance, wrote “Entanglement is a property of the quantum world; classical systems need not apply," and, “We do not endorse this new nomenclature," adding, “… this situation lacks the key feature - nonlocality" [9].
Despite the wide range of opinions, the past decade has witnessed a plethora of potentially useful experiments with vector optics and other classical implementations of vector spaces, demonstrating the power of classical entanglement (or rather non-separability) in high-precision measurements, metrology applications, and in emulation of various quantum information protocols. However, the controversy in interpreting these experiments and discussions about the fundamental nature of entanglement, both quantum and classical, remain ongoing [6]."
@@
Instead, there is a decisive difference: the characterization of quantum entanglement always involves correlating statistical outcomes of multiple measurements on a given system, as elaborated below. Conversely, classical entanglement typically refers to deterministic correlations between a single measurement and a filter or sorter operation, such as those imposed on a light beam by a polarizing beam splitter or a quantum sorter [10, 11]. The settings of such filters or sorters do not involve measurement (and is a unitary operation). This provides a clear distinction between the two phenomena.
@@ 2 Entanglement revisited: questions to ask
The non-local nature of correlations in quantum entanglement stems from non-separability. However, similar types of correlations can be found in a classical description involving mathematically non-separable functions. Are we discussing here correlations of a common nature, as suggested, for example, by Eberly? The term classical entanglement was coined. Is such terminology appropriate? Does any mathematical non-separability imply Bell-type non-locality, and is it this non-locality that makes up entanglement?"
Figure 1: Right: Different possible subsets of a set of general non-separable states. The ellipses represent the four decisive properties of systems under consideration: consist of two objects; exhibit non-locality; allows for two measurements; can be written as mathematically non-separable mode functions. The examples of physical systems populating a particular set are described in detail in the text: I - non-separable states involving two excitations (two objects); IIa,b - non-separable states with only one single excitation (object), but where two distinct measurements can still be performed on two non-separable partitions of the Hilbert space. III - classically entangled states described by non-separable mode functions. Left: Two paradigmatic examples of the states of type I and IIb, both allowing for two distinct independent measurements - see text for discussion.
@@
We want to address the problem from a different, more hands-on viewpoint. Realism, locality, and non-contextuality are all related to a measurement process. Figure 1 visually highlights some of the nontrivial aspects of partitioning and the role of measurement in analyzing correlations in non-separable systems. For simplicity, we restrict ourselves to bi-partite non-separable states. Different possible partitions of the mathematical vector space can be of quite different physical nature. They can correspond to objects of zero or non-zero rest mass, such as a photon or an atom, respectively, or to mode functions, such as particular spatial modes.
@@
To re-iterate, the complication stems from the following observation: non-separability is a mathematical property and is independent of the physical nature of the different degrees of freedom; i.e., the different parts of Hilbert space, whereas the ability to perform one or more measurements does depend on the physical nature of the degree of freedom. A measurement requires some excitation, some energy above the level of the vacuum, which can be associated either with a zero or non-zero rest mass particle or a combination thereof. Mode functions in optics, for example, can have numerous degrees of freedom occupying a larger Hilbert space, but they are merely frames for excitations and by themselves they do not refer to any excitation. If the mode is not excited, there is nothing specific to measure (other than non-specific vacuum fluctuation). We, therefore, attempt to structure the types of non-separable states with regard to their physical nature. We take the liberty here to use the word ‘excitation" in a wider sense. We speak about excitation not only when we have some non-zero energy state of the electromagnetic field, but we can also view a massive particle as a de Broglie wave-packet. We can then also think of “modes" which are “excited" or “not excited," meaning the particle is there or not. Furthermore, there is a whole range of quasi-particles, meaning a combined excitation from entities of different physical natures: phonons, excitons, polaritons, to name a few.
@@ 2.1 Defining properties of non-separable systems of different types
$|\Psi_{\rm singlet}\rangle_{12}$ $=\frac{1}{\sqrt{2}}\left(|\uparrow\rangle_{1}|\downarrow\rangle_{2}-|\downarrow\rangle_{1}|\uparrow\rangle_{2}\right),$ (1)
$|\Psi_{\rm pol}\rangle_{12}$ $=\frac{1}{\sqrt{2}}\left(|H\rangle_{1}|V\rangle_{2}-|V\rangle_{1}|H\rangle_{2}\right).$
@@
The next medium-blue shaded subset includes all states that exhibit the property of non-locality (in its trivial sense: parts of a composite system are at spatially separated locations). We briefly discuss non-locality in more general sense in Sec. 4, but it is deliberately not the subject of this paper.
@@
The larger ellipse of the yet lighter blue includes all states characterized by a common feature: the possibility to perform two distinct, independent projective measurements. Remarkably, two of the three subsets contained within the ellipse labelled ‘two measurements’ deprive us of the first traditional ingredient: two objects. Here, in these two lighter-shaded subsets, there is only one single excitation initially. However, importantly, it still clearly defines two Hilbert space partitions. These partitions are determined by two distinct excitations generated by one single object, examples in the next section. The mathematically non-separable mode functions describe the possible combination of measurement outcomes defined by these secondary excitations.
@@
For states in the outermost ellipse (but outside the other subsets), we now relinquish the possibility of performing two measurements. In this class, the only remaining property is mathematical non-separability, which is also common to all cases listed above.
@@ 2.2 Paradigmatic examples of non-separable systems
Consider a single excitation shared by two mode functions, involving non-separability. The seminal example of such quantum entanglement is the entanglement of a single photon (or atom) incident on a beam splitter. The state at the output of the beam splitter can be written as $|\psi_{12}\rangle\propto|1\rangle_{1}|0\rangle_{2}+|0\rangle_{1}|1\rangle_{2}$, where $|1\rangle_{j}$ refers to a photon (or atom) in output $j$ and $|0\rangle_{j}$ represents no photon (or atom), with $j=1,2$.
@@
In the case of an ‘undivisible’ single photon, one output mode contains a single photon while the other exhibits vacuum fluctuations in the mode defined by this excitation. Both of these secondary excitations (with one corresponding to the ground state) can be measured using homodyne detection. Thus, although only one physical excitation is present (i.e., one photon), two measurements can still be performed at the output using homodyning, providing evidence of entanglement (see also [16]). This example IIa, also demonstrates the property of non-locality (see Fig. 1).
@@
$|\Psi_{\rm 2\times 1D}\rangle_{12}$ $=\frac{1}{\sqrt{2}}\left(|n\rangle_{1}|m\rangle_{2}-|m\rangle_{1}|n\rangle_{2}\right),$ (2)
@@
Here, $\{|n,m\rangle\}$ denotes two possible energy levels associated with two excitations, either in the two oscillators $1$ and $2$, for two 1D systems, or with two spatial directions $x$ and $y$ of a 2D system.
Crucially, similar to example IIa, there are two distinct secondary excitations. Hence, despite containing only one mass, this system allows for two distinct measurements to be performed. thus decisively manifesting the entanglement property. [^1] In this example we view the quantum harmonic oscillator as a discrete-variable (DV) system due to its discrete energy levels. Alternatively, one can view it as a continuous-variable (CV) system by looking at position and momentum, which requires a different framework. Note that a light mode in quantum optics can be represented as such a CV quantum harmonic oscillator [14], and in example IIa, the two output modes can be viewed as such two coupled harmonic oscillators.
A 2D oscillator provides a prototypic example of such states, but the subset is not restricted to 2D systems. Non-separable states involving quantum entanglement between two different degrees of freedom exhibit the same property. A vivid example of locally entangled state is a mesoscopic Schorödinger cat-like state of cold atoms realized with trapped Be ions in 1996 in the group of Wineland [18]. They created the following state of a single ${}^{9}{\rm Be}^{+}$ laser-cooled ion:
@@
All the examples discussed are characterized by mathematically non-separable functions, as shown in Fig. 1. Now, consider the region of the full set outside the three inner subsets. In contrast to the previous examples, here the partition of the Hilbert space is not defined by two excitations, representing a case of classical entanglement.
For classically entangled non-separable states, the Hilbert space is spanned by the available mode functions, whereas a single existing excitation can, in principle, be assigned to (included in) any of the partitions. This assignment is guided solely by experimental convenience or our choice of which measurements to perform.
Let us illustrate this with an example. Enhancement in a number of metrology tasks has been demonstrated using a classically entangled state with correlations between the transverse spatial mode and polarization (see, e.g., [4]):
$|\Psi\rangle\propto|\psi_{01}(\mathbf{r})\rangle_{1}|H\rangle_{2}-|\psi_{10}(\mathbf{r})\rangle_{1}|V\rangle_{2}.$ (4)
As we mentioned at the beginning of our discussion on Fig. 1, quantum mechanics stems from observations, namely measurements. In the case of Eq. (4), we are dealing with the non-separability of a vector space described entirely by available mathematical mode functions connected to different degrees of freedom. Initially, there is nothing specific to measure.
@@
For Eq. (4), the partitioning initially is not connected to any excitations and the eigenvalues of the basis kets are mode-functions, possible decomposition of the initial state in some basis. As we will discuss in detail later, one of the eigenvalues is the conditional measurement outcome of one single projective measurement on a chosen partition, conditioned on the deterministic choice of the value of the other variable. That is, to unlock the physical nature of the state Eq. (4) and to connect it to observations, we rather need to view Eq. (4) differently.
@@
where $E(\mathbf{r},t)$ describes the excitation of the electromagnetic field. Put simply, we have decomposed the electromagnetic wave into two spatial modes: $\psi_{01}$ horizontally polarized and $\psi_{10}$ vertically polarized. In this system, we cannot violate realism, a cornerstone of classical physics absent in quantum mechanics. Here, we are dealing with measurement outcomes predetermined before measurement, evident in the correlations observed.
@@
To elucidate the relationship between the spatial mode and polarization, we can opt for one of the two possible scenarios. First, assume we aim to measure polarization. We assign excitation to partition 2 and conduct measurement $\hat{M}_{2}$ in the basis $\{H,V\}$, where potential outcomes are denoted ${\cal M}^{j}_{2}$, with $j=H,V$. However, to discern the correlation between spatial mode and polarization, one must first choose a specific spatial mode to examine: either $\psi_{01}(\mathbf{r})$ or $\psi_{10}(\mathbf{r})$ from partition 1. This is not a measurement but rather a unitary operation  [^2] A sorter or filter introduces a certain projection in the corresponding vector space, resembling quantum measurement. However, in a sorter, the state is projected and the complementary projection is in principle still available, such as in a polarizing beam splitter in optics. This operation can be reverted if both output beams are used. In a quantum measurement one is necessarily ‘loosing’ information which is in principle there before the measurement. The quantum measurement projects the state of the system onto an eigen state of the operator describing the measurement. A measurement often involves a sorting or filtering operation but it has to be followed by some kind of read out. The readout causes the collapse or projection of the state of the system to just one possibility. With sorting without the readout all possibilities would be still open. (in contrast to the projections in the Hilbert sub-space in previous examples being always non-unitary, projective measurements). The projection onto one of the axes in partition 1 determines the axis onto which the projection in partition 2 occurs, thus predetermining the measurement outcome on 2 before any measurement is performed: a unitary operation on 1 selecting $\psi_{01}(\mathbf{r})$ implies ${\cal M}^{H}_{2}$.

### v3

@@ 1 Introduction
This exploration has prompted a controversial discussion, the essence of which we highlight with a few selected examples. In 2011, Qian and Eberly suggested a new interpretation of light polarization based on “non-quantum entanglement," stating that “polarization is a characterization of the correlation between the vector nature and the statistical nature of the light field." Subsequent elaborations on the emerging links between quantum and classical optics [7, 8] led Eberly et al. to a strong statement: “… entanglement is a vector space property, present in any theory with a vector space framework. There is no distinction between quantum and classical entanglements, as such" [8]. This assertion has elicited equally strong opposing statements. As has been highlighted in discussions on this topic in Science [9], “Entanglement is a property of the quantum world; classical systems need not apply". Karimi and Boyd wrote, referring to "classical entanglement" term: “We do not endorse this new nomenclature," adding, “… this situation lacks the key feature - nonlocality" [9].
Despite the wide range of opinions, the past decade has witnessed a plethora of potentially useful experiments with vector optics and other classical implementations of vector spaces, demonstrating the power of ’classical entanglement’ (or rather non-separability) in high-precision measurements, metrology applications, and in emulation of various quantum information protocols. However, the controversy in interpreting these experiments and discussions about the fundamental nature of entanglement, both quantum and classical, remains ongoing [6].
@@
Instead, there is a decisive difference: the characterization of quantum entanglement always involves correlating statistical outcomes of multiple measurements on a given system, as elaborated below. Conversely, classical entanglement typically refers to deterministic correlations between a single measurement and a filter or sorter operation, such as those imposed on a light beam by a polarizing beam splitter or another quantum sorter [10, 11]. The settings of such filters or sorters do not involve measurement (and is a unitary operation). This provides a clear distinction between the two phenomena.
@@ 2 Entanglement revisited: questions to ask
The non-local nature of correlations in quantum entanglement stems from non-separability. However, similar types of correlations can be found in a classical description involving mathematically non-separable functions. Are we discussing here correlations of a common nature, as suggested, for example, by Eberly? The term ’classical entanglement’ was coined. Is such terminology appropriate? Does any mathematical non-separability imply Bell-type non-locality, and is it this non-locality that makes up entanglement?
Figure 1: Left: Different possible subsets of a set of general non-separable states. The ellipses represent the four decisive properties of systems under consideration: consist of two objects; exhibit non-locality; allows for two measurements; can be written as mathematically non-separable mode functions. The examples of physical systems populating a particular set are described in detail in the text: I - non-separable states involving two excitations (two objects); IIa,b - non-separable states with only one single excitation (object), but where two distinct measurements can still be performed on two non-separable partitions of the Hilbert space. III - classically entangled states described by non-separable mode functions. Right: Two paradigmatic examples of the states of type I and IIb, both allowing for two distinct independent measurements - see text for discussion.
@@
We want to address the problem from a different, more hands-on viewpoint. Realism, locality, and non-contextuality are all related to a measurement process. Figure 1 visually highlights some of the nontrivial aspects of partitioning and the role of measurement in analyzing correlations in non-separable systems. For simplicity, we restrict ourselves to bi-partite non-separable states. Different possible partitions of the mathematical vector space can be of quite different physical nature [15]. They can correspond to objects of zero or non-zero rest mass, such as a photon or an atom, respectively, or to mode functions, such as particular spatial modes.
The crucial role of partitioning has first been explicitly highlighted early 2000s as ambiguity and hence freedom in selecting multipartite structure for a given system. [^1] Not so much thinking about different degrees of freedom and their nature. This renders also the presence of entanglement and its degree ambiguous [16, 17]. In [16] Zanardi writes: “Clearly even the notion of entanglement is affected by some ambiguity being relative to the selected multipartite structure…The above ambiguity is removed as soon as, according to some criterion, a preferred multipartite structure is selected among the family of all possible partitions into subsystems.” These papers hence aimed at presenting a way to control the amount of achievable entanglement via the partitioning choice. Their main body is devoted to develop a corresponding mathematical set-up, based on the ideas of virtual quantum systems and compoundness notion [16], leading to a general algebraic framework where the desired quantum tensor product structure is observable induced [17]. The interested reader is referred to [16, 17], as the presented theory is beyond the scope of our work and not relevant for our discussion here. We share however one of the key realisation, stated already in these papers: “a partitioning of a given Hilbert space is induced by the experimentally accessible observables (interactions and measurements)”[17]. Or putting it differently: “the system $S$ is viewed as composed by $S_{1}$, $S_{2}$, . . . if one has some operational access (is able to “access,” “control,” “measure”) to the individual degrees of freedom of $S_{1}$, $S_{2}$[16].
@@
For the task we set here, operational distinction between quantum and classical non-separability, this crucial point should be formulated in a different way: it is not the partitioning choice but the physical nature of the given subsystems that is a key issue. This consideration becomes particularly profound in the realm of optics. To re-iterate, the complication stems from the following observation: non-separability is a mathematical property and is independent of the physical nature of the different degrees of freedom; i.e., the different parts of Hilbert space, whereas the ability to perform one or more measurements does depend on the physical nature of the degree of freedom. A measurement requires some excitation, some energy above the level of the vacuum, which can be associated either with a zero or non-zero rest mass particle or a combination thereof. Mode functions in optics, for example, can have numerous degrees of freedom occupying a larger Hilbert space, but they are merely frames for excitations and by themselves they do not refer to any excitation. If the mode is not excited, there is nothing specific to measure (other than non-specific vacuum fluctuation). We, therefore, attempt to structure the types of non-separable states with regard to their physical nature. We take the liberty here to use the word ‘excitation" in a wider sense. We speak about excitation not only when we have some non-zero energy state of the electromagnetic field, but we can also view a massive particle as a de Broglie wave-packet. We can then also think of “modes" which are “excited" or “not excited," meaning the particle is there or not. Furthermore, there is a whole range of quasi-particles, meaning a combined excitation from entities of different physical natures: phonons, excitons, polaritons, to name a few. To emphasize this aspect we are writing the wave functions also in the following form, underlining that the degrees of freedoms provide the space where excitations can live:
$|\Psi\rangle$ $=|\rm excitation\rangle_{\rm degrees~{}of~{}freedom}$ (1)
@@ 2.1 Defining properties of non-separable systems of different types
$|\Psi_{\rm singlet}\rangle_{AB}$ $=\frac{1}{\sqrt{2}}\left(|\uparrow\rangle_{A}|\downarrow\rangle_{B}-|\downarrow\rangle_{A}|\uparrow\rangle_{B}\right),$ (2)
$|\Psi_{\rm pol}\rangle_{AB}$ $=\frac{1}{\sqrt{2}}\left(|H\rangle_{A}|V\rangle_{B}-|V\rangle_{A}|H\rangle_{B}\right).$
In order to facilitate the comparison between different scenarios, we will use the notation of Eq.(1) so that the above reads in terms of ‘modes’ and excitations as:
$|\Psi_{\rm singlet}\rangle$ $=\frac{1}{\sqrt{2}}\left(|1\rangle_{\rm A,\uparrow}|1\rangle_{\rm B,\downarrow}-|1\rangle_{\rm A,\downarrow}|1\rangle_{\rm B,\uparrow}\right),$ (3)
$|\Psi_{\rm pol}\rangle$ $=\frac{1}{\sqrt{2}}\left(|1\rangle_{\rm modeA,H}|1\rangle_{\rm modeB,V}-|1\rangle_{\rm modeA,V}|1\rangle_{\rm modeB,H}\right).$
Here, for example, the mode index {$A,\uparrow$} denotes ‘a mode’ understood as a framework for excitation at location $A$ in the form of a particle in spin up (see Eq. (1) and discussion before it). In the corresponding Dirac ket, $|1\rangle_{\rm A,\uparrow}$, “$1$” means that this mode is indeed filled with this excitation, that is, particle in spin up state is present at location A. The ket $|0\rangle_{\rm A,\uparrow}$ would mean no spin-up particle at $A$. In the second equation, ‘a mode’ $\rm modeA,H$} is a spatial mode $A$ framing an excitation corresponding to a photon in $A$, polarized horizonatally, with $1$ in the ket meaning “this mode is excited”, that is, we have a horizontally polarized photon.
@@
The next medium-blue shaded subset includes all states that exhibit the property of non-locality (in its trivial sense: parts of a composite system are at spatially separated locations). This may require merely one photon [18]:
$|\Psi_{\rm pol}\rangle$ $=\frac{1}{\sqrt{2}}\left(|1\rangle_{\rm modeA}|0\rangle_{\rm modeB}-|0\rangle_{\rm modeA}|1\rangle_{\rm modeB}\right).$ (4)
We briefly discuss non-locality in more general sense in Sec. 4, but it is deliberately not the subject of this paper.
@@
The larger ellipse of the yet lighter blue includes all states characterized by a common feature: the possibility to perform two distinct, independent projective measurements. Remarkably, two of the three subsets contained within the ellipse labelled ‘two measurements’ deprive us of the first traditional ingredient: two objects. Here, in these two lighter-shaded subsets, there is only one single excitation initially. However, importantly, it still clearly defines two Hilbert space partitions. These partitions are determined by two distinct excitations generated by one single object, examples are given in the next section (see second line in Eq. (6)). The mathematically non-separable mode functions describe the possible combination of measurement outcomes defined by these secondary excitations.
@@
For states in the outermost ellipse (but outside the other subsets), we now relinquish the possibility of performing two measurements. In this class, the only remaining property is mathematical non-separability, which is also common to all cases listed above. This non-separability is not tied to excitations, and may even be demonstrated with a single coherent state excitation $|\alpha\rangle$, for example as discussed in [5] at the end of sec. 4.1.: decomposition in a particular basis followed by sorting/filtering determined by this basis can yield correlations emulating those of entangled state:
$|\Psi_{\rm non-sep}\rangle$ $=|\alpha\rangle_{\left({\rm modeA,H}\right)-\left({\rm modeB,V}\right)}.$ (5)
Here, the mode function characterized by spatio-temporal mode A and horizontal polarization is written as $\{{\rm modeA,H}\}$. For details see [5].
@@ 2.2 Paradigmatic examples of non-separable systems
Consider a single excitation shared by two mode functions, involving non-separability. The seminal example of such quantum entanglement is the entanglement of a single photon (or atom) incident on a beam splitter. The state at the output of the beam splitter can be written as $|\psi_{12}\rangle\propto|1\rangle_{1}|0\rangle_{2}+|0\rangle_{1}|1\rangle_{2}$ (see Eq. (4)), where $|1\rangle_{j}$ refers to a photon (or atom) in output $j$ and $|0\rangle_{j}$ represents no photon (or atom), with $j=1,2$.
@@
In the case of an ‘undivisible’ single photon, one output mode contains a single photon while the other exhibits vacuum fluctuations in the mode defined by this excitation and vice versa. Both of these secondary excitations (with one corresponding to the ground state) can be measured using homodyne detection. Thus, although only one physical excitation is present (i.e., one photon), two measurements can still be performed at the output using homodyning, providing evidence of entanglement (see also [19]). This example IIa, also features the property of non-locality (see Fig. 1).
@@
$|\Psi_{\rm 2\times 1D}\rangle_{AB}$ $=\frac{1}{\sqrt{2}}\left(|n\rangle_{A}|m\rangle_{B}-|m\rangle_{A}|n\rangle_{B}\right),$ (6)
@@
Here, $\{|n,m\rangle\}$ denotes two possible energy levels associated with two excitations, either in the two oscillators at locations $A$ and $B$, for two 1D systems, or with two spatial directions $x$ and $y$ of a 2D system.
Crucially, similar to example IIa, there are two distinct secondary excitations. Hence, despite containing only one mass, this system allows for two distinct measurements to be performed, thus decisively manifesting the entanglement property. [^2] In this example we view the quantum harmonic oscillator as a discrete-variable (DV) system due to its discrete energy levels. Alternatively, one can view it as a continuous-variable (CV) system by looking at position and momentum, which requires a different framework. Note that a light mode in quantum optics can be represented as such a CV quantum harmonic oscillator [14], and in example IIa, the two output modes can be viewed as such two coupled harmonic oscillators.
A 2D oscillator provides a prototypic example of such states, but the subset is not restricted to 2D systems. Non-separable states involving quantum entanglement between two different degrees of freedom exhibit the same property. A vivid example of a locally entangled state is a mesoscopic Schrödinger cat-like state of cold atoms realized with trapped Be ions in 1996 in the group of Wineland [21]. They created the following state of a single ${}^{9}{\rm Be}^{+}$ laser-cooled ion:
@@
All the examples discussed are characterized by mathematically non-separable functions, as shown in Fig. 1. Now, consider the region of the full set outside the three inner subsets. In contrast to the previous examples, here the partition of the Hilbert space is not defined by two excitations, and is thus representing a case of classical entanglement.
For classically entangled non-separable states, the Hilbert space is spanned by the available mode functions, and a single excitation, which can, in principle, be assigned to (included in) any of the partitions. This assignment is guided solely by experimental convenience or our choice of which measurements to perform.
Let us illustrate this with an example. In a number of metrology tasks, performance enhancement has been demonstrated using a classically entangled state with correlations between the transverse spatial mode and polarization. This is the situation formally displayed in Eq. (5). Often, however, the distinction between excitation and mode functions is not carved out so explicitely and one finds the following notation (see, e.g., [4]):
As we mentioned at the beginning of our discussion on Fig. 1, quantum mechanics stems from observations, namely measurements. In the case of Eq. (8), we are dealing with the non-separability of a vector space described entirely by available mathematical mode functions connected to different degrees of freedom. The excitation enabling a measurement and the non-separable mode functions as a whole form a product state. This excitation, thus, does not contribute to the entanglement.
@@
For Eq. (8), the partitioning initially is not connected to any excitations and the eigenvalues of the basis kets are mode-functions, possible decomposition of the initial state in some basis. As we will discuss in detail later, one of the eigenvalues is the conditional measurement outcome of one single projective measurement on a chosen partition, conditioned on the deterministic choice of the value of the other variable. That is, to unlock the physical nature of the state Eq. (8) and to connect it to observations, we rather need to view this equation in the spirit of Eq. (5).
@@
where $E(\mathbf{r},t)$ describes the excitation of the electromagnetic field. Put simply, we have decomposed the electromagnetic wave into two spatial modes: $\psi_{01}$ horizontally polarized and $\psi_{10}$ vertically polarized (cf Eq. (5)). Modifying the notation in this way brings us a bit closer to Eq. (1), which clearly emphasizes what is the mode (subscript) and what is the excitation (argument of the ket). In this system, we cannot violate realism, a cornerstone of classical physics absent in quantum mechanics, Here, we are dealing with measurement outcomes predetermined before measurement, evident in the correlations observed.
This is the key difference to the quantum entanglement of Eq. (7), which is also local but which displays two correlated excitations. Classical non-separability of Eqs. (5), (8), (9) is associated with only this one excitation. Exactly this difference is captured in Figures 3 and 10 - we need to compare the measurement procedures, as we elaborate in detail in the next section. The two states are deceptively similar: Both of them do not posses the non-locality property and exhibit correlations between two internal degrees of freedom. Both can be can written as a non-separable state in a composite vector space. However, for the quantum state of Eq. (7), the partitioning of Hilbert space is dictated by more than one excitation present, hence two independent measurements can be performed, one on each partition, and their statistics and correlation can be verified (see also [16, 17] and our discussion on it in Sec. 2). As we already highlighted above, the situation is drastically different in the case of classical entanglement of Eq. (8); and this is not a question of being able to perform a certain type of measurement or not, but rather a problem of having nothing specific to measure in one of the partitions.
Note that the latter statement is also relevant to differentiate between IIa and III. For IIa, the input beam splitter modes determine a vacuum mode at one of the outputs of this beam splitter alongside a photon in the other one and vice versa. Tracing out one of the output modes, the remaining output mode will be in a mixed state excited by a fraction of a photon on average. Both these correlated output modes, each with a fractional excitation, can be measured demonstrating the correlation, provided the right measurement is done, which is homodyne detection in this case [18]. In a way, in case of states in III, there is a free choice of where to assign the one and only available excitation $|{E(\mathbf{r},t)}\rangle$, as illustrated by Eq. (9). Formally, the excitation can be assigned to any one of the mode function partitions. Let us elaborate on what this implies for a measurement process.
@@
To elucidate the relationship between the spatial mode and polarization, we can opt for one of the two possible scenarios. First, assume we aim to measure polarization. We assign excitation to partition 2 and conduct measurement $\hat{M}_{2}$ in the basis $\{H,V\}$, where potential outcomes are denoted ${\cal M}^{j}_{2}$, with $j=H,V$. However, to discern the correlation between spatial mode and polarization, one must first choose a specific spatial mode to examine: either $\psi_{01}(\mathbf{r})$ or $\psi_{10}(\mathbf{r})$ from partition 1. This is not a measurement but rather a unitary operation  [^3] A sorter or filter introduces a certain projection in the corresponding vector space, resembling quantum measurement. However, in a sorter, the state is projected and the complementary projection is in principle still available, such as in a polarizing beam splitter in optics. This operation can be reverted if both output beams are used. In a quantum measurement one is necessarily ‘loosing’ information which is in principle there before the measurement. The quantum measurement projects the state of the system onto an eigen state of the operator describing the measurement. Often a sorting or filtering operation is followed by a measurement leading to some kind of read out. The readout ’causes’ the collapse or projection of the state of the system to just one possibility. With sorting only, without any readout, all possibilities would still be open. (Note, that if you were to look only at the read out in one of the output ports of a sorter, what you see is independent of what is done at the other sorter outputs. This, by the way, is the reason why entanglement does not allow one to transfer information faster than the speed of light. (in contrast, combining the projections in the Hilbert sub-space with a measurement as in previous examples, the evolution is always non-unitary, as for any projective measurements). The projection onto one of the axes in partition 1 determines the axis onto which the projection in partition 2 occurs, thus predetermining the measurement outcome on 2 before any measurement is performed: a unitary operation on 1 selecting $\psi_{01}(\mathbf{r})$ implies ${\cal M}^{H}_{2}$.